4 gm of a gas occupy 22.4 litres at into . The specific heat of gas at constant vol is 5 j/k .if speed of any quantity x of gas is at into is 952m/s. Then heat capacity at constant pressure is(R=8.3j/k/mol)
1
Expert's answer
2019-07-24T15:51:10-0400

The volume of a mole of any ideal gas at NTP (normal temperature and pressure) is 22.4 L. Since 4.0 g of a gas occupies 22.4 liters at NTP, so the molecular mass of the gas is M=4.0 g mol-1=4.0·10-3 kg mol-1.

The speed of the sound in the gas is


"v=\sqrt{\frac{\gamma RT}{M}}"

where "\gamma" is the ratio of the heat capacities at constant pressure, Cp, and at constant volume, Cv,


"\gamma =\frac{{{C}_{p}}}{{{C}_{v}}}"

R=8.3 J K-1 mol-1 is the universal gas constant, T is the temperature of the gas (T=273 K at NTP). We are also given that v=952 m s-1.

Find "\gamma" :


"{{v}^{2}}=\frac{\gamma RT}{M}\,\,\,\Rightarrow \,\,\,\gamma =\frac{M{{v}^{2}}}{RT}"


Substitute known values


"\gamma =\frac{4.0\cdot {{10}^{-3}}\text{ kg mo}{{\text{l}}^{-1}}\cdot {{\left( 952\,\text{m}\,{{\text{s}}^{-1}} \right)}^{2}}}{8.3\text{ J }{{\text{K}}^{-1}}\text{mo}{{\text{l}}^{-1}}\cdot 273\,\text{K}}=1.6"


Now find heat capacity at constant pressure


"\gamma =\frac{{{C}_{p}}}{{{C}_{v}}}\,\,\,\Rightarrow \,\,\,{{C}_{p}}=\gamma {{C}_{v}}"


Since "\gamma" =1.6 and Cv=5 J K-1mol-1, then


"{{C}_{p}}=1.6\cdot 5\text{ J }{{\text{K}}^{-1}}\text{mo}{{\text{l}}^{-1}}=8\,\text{J }{{\text{K}}^{-1}}\text{mo}{{\text{l}}^{-1}}"